OpenStudy (anonymous):

What is the volume of a square pyramid with base edges of 20 meters and a height of 30 meters?

OpenStudy (campbell_st):

V= 1/3 base area x perp height V = 1/3 20^2 x 30

OpenStudy (anonymous):

4000 m^3